SPECIATION OF PLUTONIUM IN WATER

\underline{JEAN-PHILIPPE BLAUDEAU}, Department of Chemistry, The Ohio State University, Columbus, OH 43210; STAN A. ZYGMUNT, Chemical Technology Division/Materials Science Division, Argonne National Laboratory, Argonne, Illinois 60439; BRUCE E. BURSTEN, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

The speciation of plutonium in water was modeled using both ab initio quantum chemistry and density functional theory (DFT) methods. Aqueous plutonium exists in four oxidation states: +3, +4, +5, and +6. The first two forms exist as the bare ions, but the +5 and +6 states exist as (PuO_{2})^{+} and (PuO_{2})^{2+} species, respectively. The coordination number for the number of water molecules around each of these species is determined and compared to experiment - which uses XANES (X-ray Absorption Near-Edge Spectroscpy) and eXAFS (extended X-ray Absorption Fine Structure) spectroscopy. The DFT calculations include generalized gradient corrections and the quantum chemistry calculations are at the spin-orbit configuration interaction level.